# Can Large Language Models be Good Path Planners? A Benchmark and Investigation on Spatial-temporal Reasoning

### v1

@@ Abstract
Large language models (LLMs) have achieved remarkable success across a wide spectrum of tasks; however, they still face limitations in scenarios that demand long-term planning and spatial reasoning. To facilitate this line of research, in this work, we propose a new benchmark, termed Path Planning from Natural Language (PPNL). Our benchmark evaluates LLMs’ spatial-temporal reasoning by formulating “path planning” tasks that require an LLM to navigate to target locations while avoiding obstacles and adhering to constraints. Leveraging this benchmark, we systematically investigate LLMs including GPT-4 via different few-shot prompting methodologies and BART and T5 of various sizes via fine-tuning. Our experimental results show the promise of few-shot GPT-4 in spatial reasoning, when it is prompted to reason and act interleavedly, although it still fails to make long-term temporal reasoning. In contrast, while fine-tuned LLMs achieved impressive results on in-distribution reasoning tasks, they struggled to generalize to larger environments or environments with more obstacles.
@@ 1 Introduction
Our experimental results suggest a strong capability of GPT-4 in spatial reasoning, when it is prompted effectively. Specifically, the action-and-effect prompting leads to 21.5% improvement in success rate than the naive few-shot prompting, implying that guiding LLMs to reason about the situated knowledge can help them make better immediate decisions. Prompting GPT-4 to reason step-by-step (i.e., CoT) further enhances its performance by 3%, whereas additionally allowing the model to perceive the environment signals (i.e., ReAct) yields the best success rate, 96.1%. However, we also note that this success is mostly limited to relatively “local” reasoning; for example, the ReAct-prompted GPT-4 is shown to plan only a few steps ahead and have to frequently adjust its plan based on the environment signals. In summary, this indicates that even the state-of-the-art GPT-4 still falls short in long-term, temporal reasoning. In contrast, while the fine-tuned LLMs can achieve a close-to-perfect success rate when planning in environments similar to where they were trained (i.e., in-distribution environments), they suffer from the lack of generalization and cannot perform well in environments with different grid sizes or different number of obstacles. We will release our benchmark, along with the source code to controllably synthesize the grid environment, as well as our LLM implementation, to the public for future advancement  [^1] Our source code and data are available on the following link.
@@ 6 Discussion and Analysis
Cost considerations: Table 13 provides the average cost per sample for all approaches. ReAct prompting has an average cost far exceeding other methods, as it repeatedly queries the API to adjust the path. This high cost renders it impractical for applications involving long distance navigation. Hence, it is crucial to develop less costly approaches.
@@ 8 Conclusions and Future Work
In this work, we analyze the ability of LLMs to perform spatial-temporal reasoning through using a newly proposed dataset, focusing path planning. Our findings highlight LLMs’ proficiency in spatial reasoning when it is guided by situated spatial information and continuous feedback from the environment they are grounded to. However, they face significant challenges when faced with scenarios requiring long-term planning and complex environments. While ReAct prompting achieved good results, this approach can be expensive and inapplicable to situations requiring long distance navigation, or finding the optimal path. Fine-tuned models were able to perform well on this task, however, their success is limited solely to the environments similar to those encountered during in training, making them inapplicable to a large variety of applications that require robustness and adaptation to unfamiliar settings. Finding approaches that can improve LLMs’ ability to perform path planning efficiently and optimally can, therefore, open the door to a wide array of application of LLMs in fields such as robotics and motion planning. We believe that our analysis and the proposed task can serve as a valuable resource for future work exploring this topic.
@@ B.1 Naive Prompts
{tcolorbox}
[enhanced,attach boxed title to top center=yshift=-3mm,yshifttext=-1mm, colback=black!5!white,colframe=black!75!black,colbacktitle=white!95!black, title=5-shot Prompt,fonttitle=,coltitle=black, boxed title style=size=small,colframe=black!50!black ,before upper= ]
@@
Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). Go from (0,1) to (3,4)
Actions: right right right down down down
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,5) and (1,2). Go from (5,4) to (0,5)
Actions: up up up up up right
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,5) and (5,2). Go from (4,2) to (0,5)
Actions: up up up right right up right
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,5), (4,2), (3,3) and (0,4). Go from (1,5) to (3,1)
Actions: left left left left down down
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,2) to (1,2)
Actions: up up up
{tcolorbox}
[enhanced,attach boxed title to top center=yshift=-3mm,yshifttext=-1mm, colback=black!5!white,colframe=black!75!black,colbacktitle=white!95!black, title=10-shot Prompt,fonttitle=,coltitle=black, boxed title style=size=small,colframe=black!50!black , before upper= ]
@@
Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). Go from (0,1) to (3,4)
Actions: right right right down down down
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4). Go from (5,4) to (2,4)
Actions: up up up
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4) and (1,5). Go from (0,5) to (1,1)
Actions: Goal not reachable
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,5) and (5,0). Go from (5,5) to (0,1)
Actions: up up up left up up left left left
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,5) and (5,2). Go from (4,2) to (0,5)
Actions: up up up right right up right
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,1) and (4,2). Go from (1,5) to (0,5)
Actions: up
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,4), (4,4), (5,3) and (4,5). Go from (0,4) to (5,5)
Actions: Goal not reachable
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,1), (4,4), (1,4) and (1,5). Go from (5,5) to (3,0)
Actions: up up left left left left left
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,2) to (1,2)
Actions: up up up
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,5), (5,0), (5,4), (0,0) and (5,3). Go from (5,2) to (2,4)
Actions: up up up right right
{tcolorbox}
[enhanced,attach boxed title to top center=yshift=-3mm,yshifttext=-1mm, colback=black!5!white,colframe=black!75!black,colbacktitle=white!95!black, title=15-shot Prompt,fonttitle=,coltitle=black, boxed title style=size=small,colframe=black!50!black ,before upper= ]
@@
Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. ###
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). Go from (0,1) to (3,4)
Actions: right right right down down down
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4). Go from (5,4) to (2,4)
Actions: up up up
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,3). Go from (2,4) to (4,3)
Actions: left down down
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4) and (1,5). Go from (0,5) to (1,1)
Actions: Goal not reachable
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,5) and (5,0). Go from (5,5) to (0,1)
Actions: up up up left up up left left left
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (4,3) and (2,3). Go from (5,5) to (5,4)
Actions: left
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,5) and (5,2). Go from (4,2) to (0,5)
Actions: up up up right right up right
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,1) and (4,2). Go from (1,5) to (0,5)
Actions: up
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,0), (0,0) and (1,3). Go from (3,2) to (3,1)
Actions: left
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,5), (4,2), (3,3) and (0,4). Go from (1,5) to (3,1)
Actions: left left left left down down
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,1), (4,4), (1,4) and (1,5). Go from (5,5) to (3,0)
Actions: up up left left left left left
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,4), (4,4), (5,3) and (4,5). Go from (0,4) to (5,5)
Actions: Goal not reachable.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,2) to (1,2)
Actions: up up up
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,5), (5,0), (5,4), (0,0) and (5,3). Go from (5,2) to (2,4)
Actions: up up up right right
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,0), (2,3), (1,2), (2,5) and (0,0). Go from (4,3) to (5,4)
Actions: right down
@@ B.2 Actions and Effects Prompt
{tcolorbox}
[enhanced,attach boxed title to top center=yshift=-3mm,yshifttext=-1mm, colback=black!5!white,colframe=black!75!black,colbacktitle=white!95!black, title=Actions and Effects Prompt,fonttitle=,coltitle=black, boxed title style=size=small,colframe=black!50!black ,before upper= ]
Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). Go from (0,1) to (3,4)
Actions: Go right. You are now at (0,2). Go right. You are now at (0,3). Go right. You are now at (0,4). Go down. You are now at (1,4). Go down. You are now at (2,4). Go down. You are now at (3,4). Hence, the action sequence is: right right right down down down
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,5) and (1,2). Go from (5,4) to (0,5)
Actions: Go up. You are now at (4,4). Go up. You are now at (3,4). Go up. You are now at (2,4). Go up. You are now at (1,4). Go up. You are now at (0,4). Go right. You are now at (0,5). Hence, the action sequence is: up up up up up right
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,5) and (5,2). Go from (4,2) to (0,5)
Actions: Go up. You are now at (3,2). Go up. You are now at (2,2). Go up. You are now at (1,2). Go right. You are now at (1,3). Go right. You are now at (1,4). Go up. You are now at (0,4). Go right. You are now at (0,5). Hence, the action sequence is: up up up right right up right
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,5), (4,2), (3,3) and (0,4). Go from (1,5) to (3,1)
Actions: Go left. You are now at (1,4). Go left. You are now at (1,3). Go left. You are now at (1,2). Go left. You are now at (1,1). Go down. You are now at (2,1). Go down. You are now at (3,1). Hence, the action sequence is: left left left left down down
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,2) to (1,2)
Actions: Go up. You are now at (3,2). Go up. You are now at (2,2). Go up. You are now at (1,2). Hence, the action sequence is: up up up
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4) and (1,5). Go from (0,5) to (1,1)
Actions: Goal not reachable.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,4), (4,4), (5,3) and (4,5). Go from (0,4) to (5,5)
Actions: Goal not reachable.

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable success across a wide spectrum of tasks; however, they still face limitations in scenarios that demand long-term planning and spatial reasoning. To facilitate this line of research, in this work, we propose a new benchmark, termed Path Planning from Natural Language (PPNL). Our benchmark evaluates LLMs’ spatial-temporal reasoning by formulating “path planning” tasks that require an LLM to navigate to target locations while avoiding obstacles and adhering to constraints. Leveraging this benchmark, we systematically investigate LLMs including GPT-4 via different few-shot prompting methodologies as well as BART and T5 of various sizes via fine-tuning. Our experimental results show the promise of few-shot GPT-4 in spatial reasoning, when it is prompted to reason and act interleavedly, although it still fails to perform long-term temporal reasoning. In contrast, while fine-tuned LLMs achieved impressive results on in-distribution reasoning tasks, they struggled to generalize to larger environments or environments with more obstacles.
@@ 1 Introduction
Our experimental results suggest that the GPT-4 model does possess some capability to perform spatial reasoning, when it is prompted effectively. Specifically, the action-and-effect prompting leads to 21.5% improvement in success rate than the naive few-shot prompting, implying that guiding LLMs to reason about the situated knowledge can help them make better immediate decisions. Prompting GPT-4 to reason step-by-step (i.e., CoT) further enhances its performance by 3%, whereas additionally allowing the model to perceive the environment signals (i.e., ReAct) yields the best success rate, 96.1%. However, we also note that this success is mostly limited to relatively “local” reasoning; for example, the ReAct-prompted GPT-4 is shown to plan only a few steps ahead and have to frequently adjust its plan based on the environment signals. In summary, this indicates that even the state-of-the-art GPT-4 still falls short in long-term, temporal reasoning. In contrast, while the fine-tuned LLMs can achieve a close-to-perfect success rate when planning in environments similar to where they were trained (i.e., in-distribution environments), they suffer from the lack of generalization and cannot perform well in environments with different grid sizes or different number of obstacles. We will release our benchmark, along with the source code to controllably synthesize the grid environment, as well as our LLM implementation, to the public for future advancement  [^1] Our source code and data are available on the following link..
@@ 2.2 Data Generation Process
Table 1: PPNL in Comparison to other Benchmarks
Benchmark Spatial Task Complexity Temporal Task Complexity Task Environment Synthetic vs Human anno. Optimization Needed? Vision Decoupled?
(Patel & Pavlick, 2022)
Understanding
None
2D Grid
Synthetic
No
Yes
Ruis et al. (2020)
Reasoning
Short-horizon Planning
2D Grid
Synthetic
No
No
Shridhar et al. (2020)
Reasoning
Short-horizon planning
Embodied Environment
Human-annotated
No
No
Côté et al. (2018)
Reasoning
Short-horizon planning
Embodied Environment
Synthetic
No
Yes
Shridhar et al. (2021)
Reasoning
Short-horizon planning
Embodied Environment
Human-annotated
No
Yes
Valmeekam et al. (2022)
Reasoning
Short-horizon planning
Blocksworld
Synthetic
Yes
Yes
PPNL (ours)
Reasoning
Long-horizon Planning
2D Grid
Synthetic
Yes
Yes
@@ 6 Discussion and Analysis
Can the results scale up to more realistic environments? One potential concern regarding this work is whether our findings apply to complex, real-life-like navigation scenarios. To assess this, we use GPT4-V(ision) with the Naive, AE and CoT prompting under 0-shot conditions on a 3D maze environment from Plaku et al. (2018). We observe similar trends to our experimental results.
Does good performance on PPNL imply reasoning or pattern imitation? Recent work (Chung et al., 2022; Ye et al., 2023) has shown that LLMs are good at pattern-based instruction following. An alternative interpretation of our results could, hence, be that the LLMs perform such imitation, and is limited to simpler scenarios. Nevertheless, we believe that this not the correct interpretation; path planning is an optimization problem (refer to A for more details) that cannot be solved using trivial pattern imitation. In addition, if the LLM were merely performing pattern imitation, we would expect to see similar scores for both the exact match accuracy and success rate. However, we notice a significant gap between the two, especially in the multi-goal setting (Tables 5 and 6).
LLMs do not show any mature ability for path planning. Simply looking at the IID scores (Tables 3 and 5) may lead one to think that LLMs are good enough at this task. However, we believe that the path planning capability would only be assured if a technique performs well on both IID and OOD subsets. This is, because, successful path planning ability means that the model is able to learn a strategy to solve a task, which would be independent of the path length and environment complexity. Our current experimental results indicate a gap between IID and OOD generalization, which thus implies that LLMs are still not mature in path planning.
@@ 8 Conclusions and Future Work
In this work, we analyze the ability of LLMs to perform spatial-temporal reasoning using a newly proposed dataset, called PPNL, focused on path planning. Our findings highlight LLMs’ potential in spatial reasoning when it is guided by situated spatial information and continuous feedback from the environment they are grounded to. However, they face significant challenges when faced with scenarios requiring long-term planning and complex environments. While ReAct prompting achieved good results, this approach can be expensive and inapplicable to situations requiring long-distance navigation, or finding the optimal path. Fine-tuned models were able to perform well on this task, however, their success is limited solely to environments similar to those encountered during training, making them inapplicable to a large variety of applications. Finding approaches that can improve LLMs’ ability to perform path planning efficiently and optimally can, therefore, open the door to a wide array of applications of LLMs in fields such as robotics and motion planning. We believe that our analysis and the proposed benchmark and task can serve as a valuable resource for future work exploring this topic.
@@ B.1 Naive Prompts
5-shot Prompt Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). Go from (0,1) to (3,4) Actions: right right right down down down ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,5) and (1,2). Go from (5,4) to (0,5) Actions: up up up up up right ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,5) and (5,2). Go from (4,2) to (0,5) Actions: up up up right right up right ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,5), (4,2), (3,3) and (0,4). Go from (1,5) to (3,1) Actions: left left left left down down ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,2) to (1,2) Actions: up up up
@@
10-shot Prompt Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). Go from (0,1) to (3,4) Actions: right right right down down down ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4). Go from (5,4) to (2,4) Actions: up up up ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4) and (1,5). Go from (0,5) to (1,1) Actions: Goal not reachable ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,5) and (5,0). Go from (5,5) to (0,1) Actions: up up up left up up left left left ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,5) and (5,2). Go from (4,2) to (0,5) Actions: up up up right right up right ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,1) and (4,2). Go from (1,5) to (0,5) Actions: up ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,4), (4,4), (5,3) and (4,5). Go from (0,4) to (5,5) Actions: Goal not reachable ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,1), (4,4), (1,4) and (1,5). Go from (5,5) to (3,0) Actions: up up left left left left left ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,2) to (1,2) Actions: up up up ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,5), (5,0), (5,4), (0,0) and (5,3). Go from (5,2) to (2,4) Actions: up up up right right
@@
15-shot Prompt Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). Go from (0,1) to (3,4) Actions: right right right down down down ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4). Go from (5,4) to (2,4) Actions: up up up ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,3). Go from (2,4) to (4,3) Actions: left down down ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4) and (1,5). Go from (0,5) to (1,1) Actions: Goal not reachable ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,5) and (5,0). Go from (5,5) to (0,1) Actions: up up up left up up left left left ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (4,3) and (2,3). Go from (5,5) to (5,4) Actions: left ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,5) and (5,2). Go from (4,2) to (0,5) Actions: up up up right right up right ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,1) and (4,2). Go from (1,5) to (0,5) Actions: up ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,0), (0,0) and (1,3). Go from (3,2) to (3,1) Actions: left ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,5), (4,2), (3,3) and (0,4). Go from (1,5) to (3,1) Actions: left left left left down down ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,1), (4,4), (1,4) and (1,5). Go from (5,5) to (3,0) Actions: up up left left left left left ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,4), (4,4), (5,3) and (4,5). Go from (0,4) to (5,5) Actions: Goal not reachable. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,2) to (1,2) Actions: up up up ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,5), (5,0), (5,4), (0,0) and (5,3). Go from (5,2) to (2,4) Actions: up up up right right ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,0), (2,3), (1,2), (2,5) and (0,0). Go from (4,3) to (5,4) Actions: right down
@@ B.2 Actions and Effects Prompt
Actions and Effects Prompt Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). Go from (0,1) to (3,4) Actions: Go right. You are now at (0,2). Go right. You are now at (0,3). Go right. You are now at (0,4). Go down. You are now at (1,4). Go down. You are now at (2,4). Go down. You are now at (3,4). Hence, the action sequence is: right right right down down down ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,5) and (1,2). Go from (5,4) to (0,5) Actions: Go up. You are now at (4,4). Go up. You are now at (3,4). Go up. You are now at (2,4). Go up. You are now at (1,4). Go up. You are now at (0,4). Go right. You are now at (0,5). Hence, the action sequence is: up up up up up right ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,5) and (5,2). Go from (4,2) to (0,5) Actions: Go up. You are now at (3,2). Go up. You are now at (2,2). Go up. You are now at (1,2). Go right. You are now at (1,3). Go right. You are now at (1,4). Go up. You are now at (0,4). Go right. You are now at (0,5). Hence, the action sequence is: up up up right right up right ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,5), (4,2), (3,3) and (0,4). Go from (1,5) to (3,1) Actions: Go left. You are now at (1,4). Go left. You are now at (1,3). Go left. You are now at (1,2). Go left. You are now at (1,1). Go down. You are now at (2,1). Go down. You are now at (3,1). Hence, the action sequence is: left left left left down down ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,2) to (1,2) Actions: Go up. You are now at (3,2). Go up. You are now at (2,2). Go up. You are now at (1,2). Hence, the action sequence is: up up up ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4) and (1,5). Go from (0,5) to (1,1) Actions: Goal not reachable. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,4), (4,4), (5,3) and (4,5). Go from (0,4) to (5,5) Actions: Goal not reachable.